# When AI Takes the Wheel: Security Analysis of Framework-Constrained Program Generation

### v1

@@ 4.1. Diversity of Scenarios
The first axis of our analysis explores how the diversity of scenarios impacts the security of Chrome extensions generated by LLMs. Building on the methodology from Section 3, we categorize the 140 scenarios into 10 distinct categories based on their core functionalities. These categories include:
Authentication & Identity (AUTH) scenarios focus on managing user credentials, authentication tokens, and identity verification across websites (e.g., AUTH06 for RSA key management).
@@
Content Manipulation (CM) scenarios modify webpage content, inject functionality, or enhance existing web interfaces (e.g., CM03 which adds custom email templates to Gmail). These extensions interact directly with webpages to modify their appearance, functionality, or behavior to enhance user experience.
Cookie Management (COOK) scenarios handle browser cookies, often reading, writing, or manipulating cookie data across domains. For example, COOK06 involves extraction of specific YouTube cookies by checking certain criteria, which requires careful implementation to prevent unauthorized access to sensitive cookie data from malicious websites or scripts.
@@
Download Management (DM) scenarios handle file downloads within the browser. Examples like DM10 for downloading YouTube videos directly, and DM03 for downloading crossword puzzles from the New York Times and other sources. These extensions frequently access Chrome’s download APIs, which can lead to security issues if download sources aren’t properly validated.
Developer Tools (DT) scenarios provide additional functionality for web developers such as debugging or API testing (e.g., DT02 which offers A/B testing functionality for web pages).
External Communication (EC) scenarios involve integrating extensions with external websites, services, or APIs. These extensions often require careful management of cross-origin communication and data sharing. For example, EC01 assists Amazon sellers in managing their accounts by processing API requests, which requires secure handling of authentication and cross-origin data flows.
@@
Workflow Automation (WA) scenarios automate repetitive tasks, form filling, or business processes across websites (e.g., WA07 for automating social interactions on Facebook posts, or WA11 for automated trading with buy/sell orders based on market conditions).
Table 1 presents the vulnerability rates across different scenario categories, while Figure 4 displays a detailed heatmap showing vulnerability frequency for the top five scenarios from each functional category. From Table 1, the evaluated models show consistently high vulnerability rates on Authentication & Identity and Cookie Management scenarios, with Claude-3.5-Sonnet exhibiting vulnerability rates of 83% and 78% respectively. These categories are particularly problematic as they handle sensitive user data and require proper implementation of Chrome’s privileged APIs. Similarly, the scenarios in Bookmark/History/Tab Management and External Communication categories also consistently trigger high vulnerability rates across most models, with DeepSeek-R1 generating vulnerable extensions in 74% of BHT scenarios. Figure 4 reveals that even within the same functional category, vulnerability rates can vary significantly between specific scenarios. However, certain scenarios consistently produce vulnerable extensions across all evaluated models. For example, scenarios in the Cookie Management, Download Management, and Authentication & Identity categories show uniformly high vulnerability rates (indicated by darker colors in Figure 4). Overall, the results indicate that LLMs often struggle to securely implement extensions that handle sensitive browser data or interact with external resources when prompted without explicit and professional security instructions, underscoring the need for targeted improvements in these security-critical areas.
Summary: LLMs frequently generate vulnerable Chrome extensions, especially in Authentication & Identity and Cookie Management scenarios (up to 83% and 78%, respectively), revealing that LLMs commonly struggle with security-critical functionalities involving sensitive data handling. This pattern indicates fundamental gaps in LLMs’ understanding of Chrome’s privilege boundaries when generating data-sensitive programs.
@@ 4.2. Diversity of LLMs
The second axis of our analysis examines how different LLMs perform in generating secure Chrome extensions. Table 2 shows that all evaluated LLMs produce vulnerable Chrome extensions, with varying frequency and severity. The analysis achieved high coverage with CoCo successfully analyzing 94.7%-98.8% of generated extensions across all models (i.e., except few cases with parsing errors), ensuring robust and reliable vulnerability detection. Llama-3.1-405B shows the highest vulnerability rate at 50%, meaning half of our scenarios (70 out of 140) resulted in vulnerable extensions when tested with this model. DeepSeek-R1 follows closely at 49.3%, while Gemini-2.0-Pro exhibits vulnerabilities in 44.3% of scenarios. When considering the absolute number of vulnerable extensions, DeepSeek-R1 produced the highest count (277), followed by Gemini-2.0-Pro (251). GPT-4o and Gemini-2.0-Flash-Thinking generated substantially fewer vulnerable extensions (115 and 108 respectively).
Another interesting finding is that advancing model capabilities may not necessarily translate to improved security outcomes. For example, from Gemini-1.5-Pro to Gemini-2.0-Pro, we observe an increase in both vulnerability rate (30.70% to 44.30%) and average vulnerabilities per scenario (2.74 to 4.05). Although o3-mini, Gemini-2.0-Flash-Thinking, and DeepSeek-R1 are advanced reasoning models, the results show that they potentially introduce more vulnerabilities. For example, o3-mini produced a high density of vulnerabilities (4.83 per scenario), while DeepSeek-R1 had the highest absolute number of vulnerable extensions (277). This suggests a concerning trend: models optimized for general programming capabilities or reasoning may not inherently prioritize security-aware code generation, and in some cases, more capable models might even introduce more complex vulnerability patterns when generating Chrome extensions.
Summary: All evaluated LLMs generate vulnerable Chrome extensions at high rates (18%-50%), with advanced reasoning models like DeepSeek-R1 and o3-mini producing vulnerabilities more frequently or with higher density. This suggests that the enhanced coding capabilities do not translate to improved security awareness in program generation.
@@ 4.3. Diversity of Vulnerabilities
Table 3 presents the distribution of four major vulnerability categories. Storage-related vulnerabilities (cookies, bookmarks/history, and other storage) are consistently the most common across all models, with Gemini-2.0-Pro and DeepSeek-R1 exhibiting particularly high numbers. Cross-origin requests and file download vulnerabilities appear frequently as well (e.g., GPT-4o with 19 and 13 instances respectively). Code Execution vulnerabilities, while less prevalent, remain a critical concern, particularly in models like Gemini-2.0-Pro (21 instances), highlighting the risk of dynamic code injection when LLMs overlook Chrome’s privilege boundaries. Table 4 further details the distribution of vulnerabilities across different scenarios. To gain deeper insight into specific vulnerable data flows, we focus our detailed analysis on GPT-4o-generated programs due to its widespread use and strong performance in coding benchmarks (White et al., 2025). Table 2 has revealed that GPT-4o produced 115 vulnerable extensions across 41 distinct scenarios, providing a representative sample of how even advanced models struggle with Chrome’s security architecture.
Code Execution.  Code execution vulnerabilities arise when an extension allows arbitrary JavaScript code to be dynamically executed through APIs like chrome.tabs.executeScript, fundamentally mishandling Chrome’s security model. This type of vulnerability is particularly dangerous as it enables attackers to inject and execute malicious code within the context of web pages, potentially leading to data theft, website manipulation, or other malicious activities (Lekies et al., 2015).
As shown in Table 4, code execution vulnerabilities are relatively rare compared to other vulnerability types, appearing in only 3 out of 140 scenarios (2.14%) generated by GPT-4o. Specifically, these vulnerabilities were identified in scenarios EC09 (a bookkeeping automation extension, 2 instances), WA01 (an SCM login automation tool, 1 instance), and WA10 (a job application automation extension, 1 instance). These scenarios share a common theme: they all involve automating interactions across websites, suggesting that when LLMs implement cross-site automation features, they tend to introduce unsafe dynamic code execution patterns.
@@
However, the distribution of code execution vulnerabilities shows considerable randomness across models as shown in Table 4. o3-mini produced all five of its vulnerable extensions under a single scenario (WA01), while Claude-3.5-Sonnet distributed its six vulnerabilities across three different scenarios (DH16, WA01, and WA10). DeepSeek-R1 generated six code execution vulnerabilities spread across three entirely different scenarios (COOK05, COOK07, and DM04). Although the absolute number of vulnerabilities differs, their distribution across scenarios appears inconsistent and unpredictable, reflecting a lack of uniform security awareness. This randomness suggests that while certain prompts (e.g., automating interactions across websites) may trigger such flaws, the specific manifestation depends heavily on the model and its stochastic output, posing challenges for ensuring reliable security in framework-constrained environments.
Table 4. Vulnerability Type Occurrence Across Different Extension Scenariosre
@@
Cross-Origin Requests.  Cross-Origin Request vulnerabilities occur when an extension allows arbitrary URLs to be fetched through privileged APIs like fetch or XMLHttpRequest. This vulnerability is particularly concerning because it can bypass the Same-Origin Policy (SOP), potentially exposing sensitive information from any domain or facilitating Cross-Site Request Forgery (CSRF) attacks.
The example of Listing 2 from the DM03 scenario generated by GPT-4o demonstrates this vulnerability. The scenario requires that the extension facilitates downloading crossword puzzles from the New York Times and other sources through a context menu. The generated extension listens for external messages with an action of ”fetchCrosswordData” and a URL parameter. Then it uses fetch API to request data from the provided URL without any origin validation, allowing arbitrary cross-origin requests that bypass normal browser restrictions.
@@
From Table 4, GPT-4o generated 19 cross-origin request vulnerabilities across 7 different scenarios (5.0% of all scenarios). The most affected scenarios include DM03 (5 flows), EC07 (4 flows), and DT08 (3 flows). These scenarios share a common theme of requiring network communication or API integration. The DM03 scenario requests an extension to download crossword puzzles from various sources, while EC07 facilitates interactions between XCALLY services using HTTP requests. Similarly, DT08 integrates and tests SOAP and REST services. This pattern extends to other models, particularly in scenarios involving data handling and external communication. For example, O3-mini generated all 28 of its cross-origin request vulnerabilities from scenarios DH01, EC01, EC02, and EC07. Claude-3.5-Sonnet produced 23 such vulnerabilities concentrated in DH01, DH011, EC02, EC07, and ESI03. These findings suggest that when LLMs are prompted to implement functionality requiring external network communication or data retrieval, they frequently overlook the necessity of validating origins or sanitizing URLs before making privileged cross-origin requests.
Arbitrary File Downloads.  Arbitrary file download vulnerabilities arise when an extension allows downloading files from arbitrary URLs without proper validation or user confirmation. This vulnerability can be exploited by malicious websites or attackers to trigger unwanted downloads, potentially leading to the distribution of malware or sensitive data exfiltration.
The example in Listing 3 demonstrates this vulnerability from the scenario DM13 requiring drag-and-drop functionality for web searching and image saving. The extension uses chrome.runtime.onMessage to listen for messages from content scripts. When a message with the action "saveImage" is received, it invokes chrome.downloads.download with an unvalidated message.imageUrl. This is insecure because the message originates from a content script, which operates in a lower-privilege context and can be influenced by the hosting web page. An attacker could exploit this by sending messages with malicious URLs, causing the extension to download arbitrary files to the user’s system.
@@
From Table 3 and 4, GPT-4o generated 13 arbitrary file download vulnerabilities across 3 different scenarios (2.1% of all scenarios). The DM13 scenario exhibited the highest concentration with 8 vulnerable flows, followed by DM09 with 3 flows and AUTH03 with 2 flows. These scenarios share a common theme of providing functionality to download or save content from the web. The DM13 scenario creates an extension for drag-and-drop image saving, and DM09 manages photo downloads from a kiosk system. The AUTH03 scenario, despite focusing on authentication functionality, surprisingly included the chrome.downloads permission in its manifest and implemented download capabilities without proper validation. Similarly, other models showed vulnerability patterns in download-related scenarios. For example, Gemini-2.0-Pro generated 44 arbitrary file download vulnerabilities across six download management scenarios (DM06, DM08, DM09, DM10, DM11, and DM13). This suggests a consistent failure across all LLMs to properly validate download sources or implement appropriate user confirmation prompts when generating extensions that handle web content downloads.
@@
Listings 4 and 5 present two implementations of GPT-4o under COOK09, but they potentially expose cookies to untrusted contexts. In the first example, the extension exposes the highly privileged chrome.cookies API via onMessageExternal without validating the message origin, retrieving sensitive authentication cookies and sending them back to any requester. The second example from the same model and scenario presents an even more troubling security lapse, allowing external messages to actively manipulate cookies through the chrome.cookies.set API. Similarly, as shown in Listing 7, the generated extension in AUTH05 uses window.postMessage with a wildcard origin (“ * ”) to expose data from chrome.storage.local to any webpage, effectively creating a universal backdoor to the extension’s privileged storage without any origin validation.
@@
Tables 3 and 4 show that privileged storage access vulnerabilities are the most common across all models. GPT-4o generated 102 instances across 27 scenarios, including 36 related to cookies, 49 to bookmarks/history, and 17 to other storage types. From Table 4, these vulnerabilities frequently arise in scenarios involving cookie management, bookmark/history handling, authentication and identity, and data handler management scenarios. The distribution of privileged storage access vulnerabilities is quite dispersed rather than concentrated in several specific scenarios, suggesting a systemic weakness in how LLMs handle these APIs. For example, BHT22, which “allows users to manage their bookmarks in the cloud, replacing the new tab with a custom interface,” revealed an interesting pattern where other models did not produce any vulnerabilities under this scenario, but GPT-4o generated 3 vulnerable extensions out of 12 attempts. In fact, even more advanced models evaluated in recent benchmarks (White et al., 2025), such as Claude-3.5-Sonnet, o3-mini, and DeepSeek-R1, show similarly poor or even worse handling of privileged storage APIs. For example, these three models produce a total of 233, 213, and 322 instances of privileged storage access vulnerabilities, respectively, which is significantly higher than the 102 instances generated by GPT-4o. These models consistently implemented extension functionality that exposed sensitive browser storage without proper origin validation, often sending cookie data directly to external requesters or allowing arbitrary sources to read and modify bookmarks, history, and other protected data. Despite their sophisticated capabilities in code generation, these models appear to fundamentally misunderstand the security boundaries of browser extensions, treating privileged APIs as regular functions without recognizing their security implications.
Listing 6: Example 3 of Privileged Storage Access in AUTH05
@@
Summary: LLMs generate diverse vulnerabilities including code execution, cross-origin requests, arbitrary file downloads, and privileged storage access. The most prevalent vulnerability involves exposing sensitive browser storage (cookies, history, and bookmarks) to untrusted contexts, highlighting a widespread security gap in LLM generated programs that manage Chrome’s privileged APIs.

### v2

@@ 4.1. Diversity of Scenarios
The first axis of our analysis explores how the diversity of scenarios impacts the security of Chrome extensions generated by LLMs. Following Section 3, we categorize the 140 scenarios into 10 distinct categories based on their core functionalities. These categories include:
Authentication & Identity (AUTH) scenarios manage user credentials, authentication tokens, and identity verification across websites (e.g., AUTH06 for RSA key management).
@@
Content Manipulation (CM) scenarios modify webpage content, inject functionality, or enhance existing web interfaces (e.g., CM03, which adds custom email templates to Gmail). These extensions interact directly with webpages to modify their appearance, functionality, or behavior to enhance user experience.
Cookie Management (COOK) scenarios deal with browser cookies (i.e., reading, writing, or manipulating cookie data across domains). For example, COOK06 extracts specific YouTube cookies by checking certain criteria. Without careful implementation, malicious websites or scripts could access sensitive cookie data.
@@
Download Management (DM) scenarios handle file downloads within the browser. For example, DM10 downloads YouTube videos directly, while DM03 downloads crossword puzzles from the New York Times and other sites. These extensions rely on Chrome’s download APIs. Without proper validation of download sources, this can create security vulnerabilities.
Developer Tools (DT) scenarios provide additional functionality for web developers, such as debugging or API testing (e.g., DT02, which offers A/B testing functionality for web pages).
External Communication (EC) scenarios connect extensions with external websites, services, or APIs. For these extensions, developers must carefully handle cross-origin communication and data sharing. For example, EC01 helps Amazon sellers manage their accounts by processing API requests, which requires secure handling of authentication and cross-origin data flows.
@@
Workflow Automation (WA) scenarios handle automation of repetitive tasks across websites, such as form filling or business processes (e.g., WA07 automates Facebook post interactions, WA11 automates trading with buy/sell orders based on market conditions).
Table 1 shows vulnerability rates across scenario categories, and Figure 4 provides a heatmap of vulnerability frequency for the top five scenarios in each category. From Table 1, the evaluated models show consistently high vulnerability rates on Authentication & Identity and Cookie Management scenarios, with Claude-3.5-Sonnet exhibiting vulnerability rates of 83% and 78% respectively. These categories are particularly problematic as they handle sensitive user data and require proper implementation of Chrome’s privileged APIs. Similarly, the scenarios in Bookmark/History/Tab Management and External Communication categories also consistently trigger high vulnerability rates across most models, with DeepSeek-R1 generating vulnerable extensions in 74% of BHT scenarios. Figure 4 shows that vulnerability rates vary significantly between specific scenarios, even within the same functional category. However, some scenarios consistently result in vulnerable programs from every model we tested. For example, scenarios in the Cookie Management, Download Management, and Authentication & Identity categories show uniformly high vulnerability rates (indicated by darker colors in Figure 4). Overall, the results show that LLMs often struggle to securely implement extensions that handle sensitive browser data or interact with external resources, especially when prompted without explicit security instructions.
Summary: LLMs frequently generate vulnerable Chrome extensions, especially in Authentication & Identity and Cookie Management scenarios (up to 83% and 78%, respectively). This highlights fundamental gaps in LLMs’ understanding of Chrome’s privilege boundaries for handling sensitive data.
@@ 4.2. Diversity of LLMs
The second axis of our analysis examines how different LLMs perform in generating secure Chrome extensions. Table 2 shows that all evaluated LLMs produce vulnerable Chrome extensions, with varying frequency and severity. The analysis achieved high coverage with CoCo successfully analyzing 94.7%-98.8% of generated extensions across all models (i.e., except for a few cases with parsing errors), providing robust and reliable vulnerability detection. Llama-3.1-405B shows the highest vulnerability rate at 50%, meaning half of our scenarios (70 out of 140) resulted in vulnerable extensions when tested with this model. DeepSeek-R1 follows at 49.3%, and Gemini-2.0-Pro shows vulnerabilities in 44.3% of scenarios. In terms of absolute numbers, DeepSeek-R1 produced the highest number of vulnerable extensions (277), followed by Gemini-2.0-Pro (251). GPT-4o and Gemini-2.0-Flash-Thinking generated substantially fewer vulnerable extensions (115 and 108 respectively).
We also observe that more advanced model capabilities do not always mean better security. For example, from Gemini-1.5-Pro to Gemini-2.0-Pro, we observe an increase in both vulnerability rate (30.70% to 44.30%) and average vulnerabilities per scenario (2.74 to 4.05). Although o3-mini, Gemini-2.0-Flash-Thinking, and DeepSeek-R1 are advanced reasoning models, the results show that they potentially introduce more vulnerabilities. For example, o3-mini produced a high density of vulnerabilities (4.83 per scenario), while DeepSeek-R1 had the highest absolute number of vulnerable extensions (277). This suggests a concerning trend. Models optimized for general programming capabilities or reasoning may not inherently prioritize security-aware code generation. In fact, in some cases, more capable models might even introduce more complex vulnerability patterns when generating Chrome extensions.
Summary: All evaluated LLMs generate vulnerable Chrome extensions, with high vulnerability rates ranging from 18% to 50%. We found that advanced reasoning models like DeepSeek-R1 and o3-mini produced vulnerabilities more frequently or with higher density. This finding suggests that better coding capabilities don’t lead to more secure framework-constrained program generation.
@@ 4.3. Diversity of Vulnerabilities
Table 3 presents the distribution of four major vulnerability categories. We can see that storage-related vulnerabilities (cookies, bookmarks/history, and other storage) are the most common across all models. Cross-origin requests and file download vulnerabilities appear frequently as well (e.g., GPT-4o with 19 and 13 instances respectively). Code Execution vulnerabilities appear less frequently but pose significant risks. For example, Gemini-2.0-Pro generates 21 instances. Table 4 further details the distribution of vulnerabilities across different scenarios. To gain deeper insight into specific vulnerable data flows, we focus our detailed analysis on GPT-4o-generated programs due to its widespread use and strong performance in coding benchmarks (White et al., 2025). Table 2 reveals that GPT-4o produced 115 vulnerable extensions across 41 distinct scenarios, providing a representative sample of how even advanced models struggle with Chrome’s security architecture.
Code Execution.  Code execution vulnerabilities arise when extensions allow arbitrary JavaScript to execute dynamically through APIs like chrome.tabs.executeScript. This vulnerability is quite dangerous since it enables attackers to inject and execute malicious code within the context of web pages. This could lead to data theft, website manipulation, or other malicious activities (Lekies et al., 2015).
Table 4 shows that code execution vulnerabilities appear less frequently compared to other vulnerability types. GPT-4o produces them in only 3 of 140 scenarios (2.14%). These vulnerabilities were identified in scenarios EC09 (a bookkeeping automation extension, 2 instances), WA01 (an SCM login automation tool, 1 instance), and WA10 (a job application automation extension, 1 instance). All three scenarios involve automating interactions across websites. This pattern suggests that LLMs struggle to implement cross-site automation safely, introducing unsafe dynamic code execution.
@@
However, the distribution of code execution vulnerabilities shows considerable randomness across models as shown in Table 4. o3-mini produced all five of its vulnerable extensions under a single scenario (WA01), while Claude-3.5-Sonnet distributed its six vulnerabilities across three different scenarios (DH16, WA01, and WA10). DeepSeek-R1 generated six code execution vulnerabilities spread across three entirely different scenarios (COOK05, COOK07, and DM04). Although the absolute number of vulnerabilities differs, their distribution across scenarios appears inconsistent and unpredictable, reflecting a lack of uniform security awareness. The randomness shows that while some prompts (e.g., automating interactions across websites) trigger these vulnerabilities, their specific form varies by model and its stochastic behavior. It poses challenges for ensuring reliable security in framework-constrained program generation.
Table 4. Vulnerability Type Occurrence Across Different Extension Scenarios
@@
Cross-Origin Requests.  Cross-Origin Request vulnerabilities occur when an extension allows arbitrary URLs to be fetched through privileged APIs like fetch or XMLHttpRequest. This vulnerability is dangerous: it bypasses the Same-Origin Policy (SOP), potentially exposing sensitive information from any domain or enabling Cross-Site Request Forgery (CSRF) attacks.
The example in Listing 2 from the DM03 scenario generated by GPT-4o demonstrates this vulnerability. The scenario requires the extension to download crossword puzzles from the New York Times and other sources through a context menu. The generated extension listens for external messages with an action of ”fetchCrosswordData” and a URL parameter. Then it uses fetch API to request data from the provided URL without any origin validation. This allows arbitrary cross-origin requests that bypass normal browser restrictions.
@@
From Table 4, GPT-4o generated 19 cross-origin request vulnerabilities across 7 different scenarios (5.0% of all scenarios). The top three scenarios include DM03 (5 flows), EC07 (4 flows), and DT08 (3 flows). These scenarios all require network communication or API integration. The DM03 scenario is for downloading crossword puzzles from various sources. EC07 handles interactions between XCALLY services using HTTP requests. Similarly, DT08 integrates and tests SOAP and REST services. Other models show the same pattern, particularly in scenarios involving data handling and external communication. For example, o3-mini generated all 28 of its cross-origin request vulnerabilities from scenarios DH01, EC01, EC02, and EC07. Claude-3.5-Sonnet produced 23 such vulnerabilities, mainly in DH01, DH011, EC02, EC07, and ESI03. These findings suggest that when LLMs are asked to implement features needing external network communication or data retrieval, they often fail to validate origins or sanitize URLs. This leads them to making privileged cross-origin requests insecurely.
Arbitrary File Downloads.  Arbitrary file download vulnerabilities arise when an extension allows downloading files from arbitrary URLs without proper validation or user confirmation. Malicious websites or attackers can exploit this vulnerability to trigger unwanted downloads. This could lead to the distribution of malware or sensitive data exfiltration.
Listing 3 shows this vulnerability in scenario DM13, which requires drag-and-drop functionality for web searching and image saving. The extension uses chrome.runtime.onMessage to listen for messages from content scripts. When a message with the action "saveImage" is received, it invokes chrome.downloads.download with an unvalidated message.imageUrl. This is insecure because the message originates from a content script. Content scripts operate in a lower-privilege context and can be influenced by the hosting web page. An attacker could exploit this by sending messages with malicious URLs. This would cause the extension to download arbitrary files to the user’s system.
@@
From Table 3 and 4, GPT-4o generated 13 arbitrary file download vulnerabilities across 3 different scenarios (2.1% of all scenarios). DM13 had the most vulnerable flows (8), followed by DM09 (3) and AUTH03 (2). These scenarios share a common theme of providing functionality to download or save content from the web. The DM13 scenario creates an extension for drag-and-drop image saving, and DM09 manages photo downloads from a kiosk system. Interestingly, the AUTH03 scenario focuses on authentication functionality but still includes the chrome.downloads permission. The extension implemented download capabilities without proper validation. Similarly, other models showed vulnerability patterns in download-related scenarios. For example, Gemini-2.0-Pro generated 44 arbitrary file download vulnerabilities across six download management scenarios (DM06, DM08, DM09, DM10, DM11, and DM13). This suggests a consistent failure across all LLMs. When generating extensions that handle web content downloads, they often fail to properly validate download sources or implement appropriate user confirmation prompts.
@@
Listings 4 and 5 present two implementations by GPT-4o for COOK09, but they potentially expose cookies to untrusted contexts. In the first example, the extension exposes the highly privileged chrome.cookies API via onMessageExternal without validating the message origin, retrieving sensitive authentication cookies and sending them back to any requester. The second example from the same model and scenario shows an even more serious flaw. It allows external messages to actively manipulate cookies through the chrome.cookies.set API. Similarly, Listing 6 shows the generated extension in AUTH05 using window.postMessage with a wildcard origin (“ * ”). This action exposes data from chrome.storage.local to any webpage, creating a universal backdoor to the extension’s privileged storage without any origin validation.
@@
Tables 3 and 4 show that privileged storage access vulnerabilities are the most common across all models. GPT-4o generated 102 instances across 27 scenarios: 36 for cookies, 49 for bookmarks/history, and 17 for other storage types. From Table 4, these vulnerabilities frequently arise in scenarios involving cookie management, bookmark/history handling, authentication and identity, and data handler management scenarios. The vulnerabilities don’t concentrate in a few scenarios but spread widely, indicating that LLMs have fundamental problems handling these APIs. For example, BHT22, which “allows users to manage their bookmarks in the cloud, replacing the new tab with a custom interface,” revealed an interesting pattern where other models did not produce any vulnerabilities under this scenario, but GPT-4o generated 3 vulnerable extensions out of 12 attempts. In fact, even more advanced models evaluated in recent benchmarks (White et al., 2025), such as Claude-3.5-Sonnet, o3-mini, and DeepSeek-R1, show similarly poor or even worse handling of privileged storage APIs. For example, these three models produced 233, 213, and 322 instances respectively of privileged storage access vulnerabilities, which is considerably higher than the 102 instances generated by GPT-4o. These models consistently implemented extension functionality that exposed sensitive browser storage without proper origin validation. They often sent cookie data directly to external requesters or allowed arbitrary sources to read and modify bookmarks, history, and other protected data. Even with sophisticated coding abilities, these models appear to fundamentally misunderstand the security boundaries of browser extensions. They treat privileged APIs like regular functions, ignoring their security implications.
Listing 6: Example of Privileged Storage Access in AUTH05
@@
Summary: LLMs generate diverse vulnerabilities, including code execution, cross-origin requests, arbitrary file downloads, and privileged storage access. The most common vulnerability is exposing sensitive browser storage (cookies, history, and bookmarks) to untrusted contexts. This highlights a widespread security gap in LLM-generated programs that manage Chrome’s privileged APIs.